# Proximal Supervised Fine-Tuning

### v1

@@ 4.2 Exploration on the Potential of Models in the RL stage
Practical LLMs often conduct RL after SFT in post-training. The supervised fine-tuned models should function well as an appropriate start point of the following RL stage, avoid both overfitting and underfitting, and better stimulate the potential of RL. In this subsection, we evaluate the power of PSFT in the RL stage.
Setup.
We adopt all techniques from DAPO (Yu et al., 2025), setting clip-higher to 0.28. The RL training uses the DAPO-MATH-17k dataset (Shao et al., 2024), with detailed training configurations provided in Appendix A.1.2. We further select the checkpoints based on the highest in-domain performance for evaluation.
@@ 4.2.1 Training Dynamics on RL
Analysis 1: PSFT leaves much room for RL optimization.
The RL entropy evolution is shown in Figure 4. As low entropy limits the model’s exploration, it is notable that using PSFT as a cold start results in higher entropy throughout training, with a steeper increase.
Analysis 2: PSFT shows slow start and rapid catch-up in RL.
The in-domain performance evolution is presented in Figure 5. Thanks to its high entropy, which promotes exploration, using PSFT as the cold start leads to a relatively steep performance improvement and ultimately surpasses the performance achieved when using SFT as the cold start.
@@ 4.2.2 Detailed Evaluation
Analysis 3: PSFT achieves better in-domain performance after RL.
Table 3 presents the results of RL experiments on in-domain benchmarks. Overall, RL improves the performance of all models. A comprehensive evaluation of the in-domain tasks shows that, although PSFT initially lags behind SFT, it holds significant potential that can be further unlocked through RL.
| Method | AIME24 | AIME25 | AMC | MATH-500 | OlympidBench | Minerva | Avg. |
| Qwen2.5-7B-Instruct | |||||||
| SFT | 22.08 | 23.02 | 62.73 | 84.10 | 52.35 | 43.66 | 47.99 |
| GRPO | 27.40 | 26.15 | 70.86 | 86.60 | 58.09 | 45.27 | 52.40 |
[…]
| Llama3.1-8B-Instruct | |||||||
| SFT | 10.63 | 16.77 | 47.19 | 72.60 | 41.43 | 32.17 | 36.80 |
| GRPO | 14.06 | 17.50 | 54.38 | 77.03 | 47.13 | 33.09 | 40.53 |
[…]
| Method | GPQA | ARC-C | TruthfulQA | MMLU-Pro | SuperGPQA | HeadQA | IFEval | Avg. |
| Qwen2.5-7B-Instruct | ||||||||
| SFT | 32.89 | 92.22 | 63.14 | 58.98 | 29.02 | 74.65 | 54.42 | 57.90 |
| GRPO | 39.39 | 92.25 | 61.93 | 62.61 | 33.80 | 75.42 | 53.89 | 59.90 |
[…]
| Llama3.1-8B-Instruct | ||||||||
| SFT | 19.38 | 87.73 | 67.08 | 50.29 | 21.95 | 73.52 | 33.45 | 50.49 |
| GRPO | 31.06 | 88.57 | 64.82 | 54.39 | 27.80 | 73.81 | 31.85 | 53.19 |
[…]
Analysis 4: PSFT outperforms SFT on out-of-domain RL tasks by a large margin.
Table 4 presents the results of RL experiments on out-of-domain benchmarks. It is evident that if the original model exhibits weak capabilities during the cold-start phase, the subsequent RL training is constrained by these limitations. For instance, the IFEval results very clearly reflect this point. This further underscores the necessity of enhancing the current SFT algorithm.
@@ 4.3 Further Experiments on Human Alignment
In this section, to demonstrate the universality of PSFT, we extend our experiments to human alignment datasets, employing different base models and algorithms such as DPO (Rafailov et al., 2023). Our results show that PSFT effectively reduces the alignment tax (i.e., the generalization gap) while still leaving room for further optimization.
Setup.
We adopt a completely different setup to demonstrate the universality of PSFT further. (1) Model and Dataset. We fine-tune the pre-trained Qwen3-4B-Base (Yang et al., 2025) model with the UltraFeedback dataset (Cui et al., 2023). (2) Training algorithm. We select the chosen region of the dataset and the SFT/PSFT model on it. And for PSFT, we train a double step named PSFT. Then, use the DPO algorithm to enable the model to learn the contrastive reward signal. (3) Evaluation. Since short CoT models can not solve the complex problems, we use the ARC, GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2020), GPQA, and TruthfulQA to evaluate the alignment tax. We evaluate our models on three alignment benchmarks: MT-Bench (Zheng et al., 2024a), AlpacaEval (Dubois et al., 2024), and Arena-Hard (Li et al., 2024a). We use Qwen3-30B-A3-Instruct-2507 (Yang et al., 2025) as the judge model to provide alignment evaluation. We use llm-eval-harness (Gao et al., 2024) to evaluate the alignment tax performance. More experimental setup, see Appendix A.1.3.
4.3.1 Training Dynamics
Analysis 1: PSFT still reliably prevents entropy collapse for other domains.
As shown in Figure 6(a), the entropy loss of SFT still shows a severe sawtooth shape, indicating a potential overfitting phenomenon, while PSFT overcomes this point. Therefore, SFT leads to a severe alignment tax (Figure 7) and restricts further optimization (Table 5).
4.3.2 Human Alignment Performance
Analysis2.
Table 5 presents the results of DPO experiments on alignment benchmarks. We observe that PSFT achieves a similar effect to SFT training while avoiding entropy collapse, and it performs even better during DPO training. Moreover, using PSFT as the cold-start provides additional benefits that are further unlocked by DPO. This is consistent as reported by Xiao (2024) that when an SFT model collapses—i.e., becomes overly biased toward certain outputs with near certainty (Figure 6(b))—it can subsequently induce preference collapse in the alignment process.
Analysis 3: PSFT training exploits both positive and negative samples well.
As shown in Figure 6(b), using SFT or PSFT as the cold-start point can influence the subsequent DPO phase. PSFT and PSFT yield similar rewards on positive samples; the main difference is that prolonged training reduces the occurrence of negative samples. In contrast, SFT training results in positive samples appearing frequently while negative samples are rare, which limits the model’s ability to learn from the reward signal and instead biases it toward the human-selected data.
@@ 4.3.2 Human Alignment Performance
[…]
Analysis4.
Figure 7 shows the alignment tax results with various tasks (e.g., GSM8K, MMLU, GPQA) as the out-of-domain tasks. It indicates that: (a) PSFT and PSFT+DPO could largely maintain the general ability during the SFT stage compared to the original SFT. (b) PSFT-Prolong, even being degraded by possible overfitting issues, achieves relatively good results on alignment tax, implying the robustness of PSFT in practical scenarios.
@@ 5.1 Clipped Tokens in PSFT
Figure 8 shows the representative clipped tokens during our PSFT’s training process. As we can see, the clipped token mostly concentrates on uncertain words like “wait”, “alternatively”, and similar words that reflect certain “long thinking patterns”, which are also supposed to be learned in the following RL stage. With the training in progress, these token-clip weights are more pronounced, while other tokens become smaller. The “thinking pattern” has been gradually and smoothly learned into models via PSFT, without much disturbance to the general capabilities.
@@ 5.2 Parameter Analysis on Clipped Value
As shown in Figure 9, PSFT without clipping maintains higher entropy but suffers from large, unstable gradient norms and highly fluctuating downstream results. By introducing the trust-region clipping mechanism, PSFT strikes a better balance: it prevents entropy collapse, stabilizes gradient updates, and achieves steady improvements in performance. Across different clipping thresholds, with medium values of (e.g., 0.28) showing particularly favorable trade-offs between stability and performance.
@@ 6 Related Work
SFT is typically the initial step in the post-training pipeline. Addressing potential issues at the SFT stage can significantly enhance the robustness of the resulting further fine-tuned model. Several studies have noted that the standard cross-entropy (CE) loss may not be the most suitable objective for SFT (Li et al., 2024b; Xiao, 2024), arguing that training with CE at this stage tends to encourage the model to memorize the training data rather than acquire more generalizable capabilities. Li et al. (2024b) proposes the GEM, a game-theoretic SFT algorithm with entropy regularization to preserve diversity and mitigate forgetting. The concurrent work is Importance-weighted SFT (iw-SFT) (Qin & Springenberg, 2025) and DFT Wu et al. (2025). The key insight of iw-SFT is to reinterpret standard SFT as optimizing a loose lower bound on the RL objective in a sparse reward setting. While vanilla SFT simply maximizes the likelihood of filtered “successful” trajectories, this lower bound becomes increasingly loose as the model distribution drifts away from the reference policy. iw-SFT addresses this gap by introducing importance re-weighting, which assigns higher weights to more preferred trajectories and thus tightens the bound toward the true RL objective. DFT views SFT as a flawed policy gradient and introduces a simple probability-based reweighting that improves generalization.
Reinforcement Learning.
@@
Following SFT, RL is employed to optimize reward signals directly. Policy gradient methods (Sutton et al., 2000) provide the foundation but are prone to instability and high variance. To address this, TRPO (Schulman et al., 2015) introduces a trust-region constraint, limiting each policy update to a small KL-divergence neighborhood to ensure stable improvement. PPO (Schulman et al., 2017) further simplifies this approach with a clipped surrogate objective, striking a balance between stability and efficiency.
7 Conclusion
@@
Motivated by TRPO and PPO, we propose Proximal Supervised Fine-Tuning (PSFT). PSFT preserves a smooth entropy curve while achieving performance comparable to standard SFT at similar entropy levels, and it exhibits strong generalization. Furthermore, using PSFT as a cold-start model facilitates more effective post-training techniques such as RL optimization and DPO. Overall, PSFT presents a promising alternative to traditional SFT.
Appendix A Appendix
A.1 Experimental Details

### v2

@@ 4.2 Exploration on the Potential of Models in the RL stage
Practical LLMs often conduct RL after SFT in the post-training phase. The SFT models should serve as an appropriate starting point for the subsequent RL stage, avoiding both overfitting and underfitting, and thereby better stimulate the potential of RL. We evaluate the power of PSFT in the RL stage.
Setup. We adopt DAPO (Yu et al., 2025) with a clip-higher value of 0.28, a stable variant of GPPO. The RL training uses the DAPO-MATH-17k dataset, with detailed training configurations provided in Appendix C.7.2. We select the checkpoints based on the highest in-domain performance for evaluation.
@@ 4.2.1 Training Dynamics on RL
Finding 1: PSFT leaves much room for RL optimization. The RL entropy evolution is shown in Figure 5(a). As low entropy limits the model’s exploration, it is notable that using PSFT as a cold start results in higher entropy throughout training, with a steeper increase.
Finding 2: PSFT shows slow start and rapid catch-up in RL. The in-domain performance evolution is presented in Figure 5(b). Thanks to its high entropy, which promotes exploration, using PSFT as the cold start leads to a relatively steep performance improvement and ultimately surpasses the performance achieved when using the standard SFT as the cold start.
@@ 4.2.2 Detailed Evaluation
Finding 3: PSFT achieves better in-domain performance after RL. Table 2 presents the results of RL experiments on in-domain benchmarks. Overall, RL improves the performance of all models. A comprehensive evaluation of six in-domain tasks shows that, although PSFT initially lags behind SFT, it holds significant potential that can be further unlocked through RL.
| Qwen2.5-7B-Instruct | Llama3.1-8B-Instruct | ||||||||
| Benchmark | SFT | SFT GRPO | PSFT | PSFT GRPO | SFT | SFT GRPO | PSFT | PSFT GRPO | |
| In-Domain | AIME-24 | 22.08 | 27.40 | 19.38 | 28.13 | 10.63 | 14.06 | 10.31 | 14.27 |
| AIME-25 | 23.02 | 26.15 | 21.98 | 27.19 | 16.77 | 17.50 | 14.48 | 18.02 | |
| AMC | 62.73 | 70.86 | 62.34 | 71.72 | 47.19 | 54.38 | 46.80 | 55.23 | |
| MATH-500 | 84.10 | 86.60 | 83.35 | 87.48 | 72.60 | 77.03 | 71.98 | 77.78 | |
| OlympidBench | 52.35 | 58.09 | 51.50 | 58.50 | 41.43 | 47.13 | 39.61 | 47.74 | |
| Minerva | 43.66 | 45.27 | 43.33 | 46.83 | 32.17 | 33.09 | 32.40 | 37.55 | |
| Avg. | 47.99 | 52.40 | 46.98 | 53.31 | 36.80 | 40.53 | 35.93 | 41.77 | |
| Out-of-Domain | GPQA | 32.89 | 39.39 | 33.21 | 43.43 | 19.38 | 31.06 | 26.89 | 36.23 |
| ARC-C | 92.22 | 92.25 | 92.29 | 92.42 | 87.73 | 88.57 | 89.11 | 90.92 | |
| TruthfulQA | 63.14 | 61.93 | 67.16 | 64.84 | 67.08 | 64.82 | 68.69 | 65.24 | |
| MMLU-Pro | 58.98 | 62.61 | 59.18 | 63.65 | 50.29 | 54.39 | 56.58 | 60.92 | |
| SuperGPQA | 29.02 | 33.80 | 28.10 | 34.51 | 21.95 | 27.80 | 25.85 | 31.30 | |
| HeadQA | 74.65 | 75.42 | 75.82 | 75.89 | 73.52 | 73.81 | 77.90 | 78.12 | |
| IFEval | 54.42 | 53.89 | 73.03 | 73.73 | 33.45 | 31.85 | 69.75 | 71.02 | |
| Avg. | 57.90 | 59.90 | 61.26 | 64.06 | 50.49 | 53.19 | 59.25 | 61.96 | |
Finding 4: PSFT still outperforms the standard SFT on out-of-domain tasks by a large margin after RL. Table 2 also presents the results of RL experiments on out-of-domain benchmarks. It is evident that if the model exhibits weak capabilities during the cold-start phase, the subsequent RL training is constrained by these limitations (e.g., the results of IFEval). PSFT functions well on different base models. It further underscores the necessity of enhancing the current SFT algorithm.
@@ 4.3 Further Experiments on Human Alignment
In this section, we demonstrate the universality of PSFT by extending our experiments to human alignment datasets, employing various base models and algorithms, such as DPO (Rafailov et al., 2023).
Our results show that PSFT effectively reduces the alignment tax (i.e., the generalization gap) while still leaving room for further optimization.
Setup. We adopt a completely different setup to demonstrate the universality of PSFT further. (1) Model and Dataset. We fine-tune the pre-trained Qwen3-4B-Base (Yang et al., 2025) model with the UltraFeedback dataset (Cui et al., 2023). (2) Training pipeline. We select the chosen region of the dataset and then perform SFT/PSFT. And for PSFT, we additionally train the double steps named PSFT. Then, use the DPO algorithm to enable the model to learn the contrastive reward signal. (3) Evaluation. See Appendix C.7.3.
Finding 1: PSFT reliably prevents entropy collapse and enables DPO for alignment. As shown in Figure 6(a), the entropy loss of SFT still shows a severe sawtooth shape, indicating a potential overfitting phenomenon, while PSFT overcomes this issue. Therefore, SFT leads to a severe alignment tax (Figure 7) and restricts further optimization. Table 3 presents the results of DPO experiments on alignment benchmarks. We observe that PSFT achieves a similar effect to SFT while avoiding entropy collapse, and it performs even better during DPO training. PSFT provides a strong starting point, with DPO training unlocking further gains.
@@
| PSFT | 11.95 | 8.37 | 17.40 | 7.41 | 5.84 |
| DPO | 19.26 | 15.17 | 30.20 | 7.63 | 6.74 |
| PSFT | 11.79 | 7.49 | 11.40 | 6.83 | 4.86 |
| DPO | 23.29 | 20.13 | 36.40 | 8.51 | 6.95 |
Finding 2: PSFT training exploits both positive and negative samples well. As shown in Figure 6(b), the cold-start point can influence the subsequent DPO phase. PSFT and PSFT yield similar rewards on positive samples; the main difference is that prolonged training reduces the occurrence of negative samples. In contrast, SFT training results in positive samples appearing frequently while negative samples are rare, which limits the model’s ability to learn from the reward signal and instead biases it toward the human-selected data. This is consistent as reported by Xiao (2024) that when an SFT model becomes overly biased toward certain outputs with near certainty, it can subsequently induce preference collapse in the alignment process.
Finding 3: PSFT reduces the alignment tax. Figure 7 shows the alignment tax results with various tasks (e.g., GSM8K, MMLU, GPQA) as the out-of-domain tasks. It indicates that: (a) PSFT and PSFT+DPO could largely maintain the general ability during the SFT stage compared to the standard SFT. (b) PSFT-Prolong, even being degraded by possible overfitting issues, achieves relatively good results on alignment tax, implying the robustness of PSFT in practical scenarios.
4.4 Further Explorations of PSFT on Vision Language Models
Setup. (1) Model and Datasets. We employ the Qwen2.5-7B-VL-Instruct (Bai et al., 2025) vision language model (VLM). We utilize the text-only OpenR1-Math-4096 dataset (Face, 2025) and the Geometry-3k (Lu et al., 2021) dataset, which consists of text-image geometry problems. (2) Training Pipeline. We conduct SFT/PSFT on the VLM using the text-only dataset to inject the long CoT reasoning ability. After that, we perform RL training on the text-image pair dataset to further enhance the visual reasoning ability. (3) Evaluation. Motivation and additional experimental details are provided in Appendix C.7.4.
Finding 1: PSFT functions well on multimodal scenarios. Based on Figure 9, PSFT still effectively avoids entropy collapse while achieving comparable in-domain performance. By maintaining sufficient entropy space, the model maintains excellent performance throughout the RL process. It suggests that incorporating text-only long CoT data can enhance the reasoning ability of VLMs.
| Benchmark | Base | SFT | PSFT |
| AIME-24 | 6.15 | 8.54 | 9.27 |
| AIME-25 | 1.88 | 10.21 | 10.73 |
| AMC | 34.77 | 44.69 | 44.84 |
| MATH-500 | 65.65 | 71.70 | 70.58 |
| OlympidBench | 29.44 | 37.50 | 36.02 |
| Minerva | 30.93 | 31.85 | 32.26 |
| MMLU | 63.10 | 45.39 | 62.44 |
| MMLU-Pro | 47.90 | 40.67 | 50.55 |
| Benchmark | Geo3K | MathVerse | MATHVista | MMMU | MMMU-Pro |
| Base | 29.28 | 37.03 | 66.80 | 41.89 | 29.59 |
| GRPO | 49.42 | 42.44 | 71.10 | 40.00 | 28.61 |
| SFT | 33.78 | 38.83 | 68.00 | 40.78 | 23.58 |
| GRPO | 48.92 | 43.12 | 72.70 | 38.78 | 19.60 |
| PSFT | 33.61 | 38.25 | 69.60 | 43.33 | 27.63 |
| GRPO | 50.58 | 44.14 | 72.20 | 40.78 | 26.07 |
Finding 2: PSFT preserves the general ability, while SFT compromises. Table 5 shows that PSFT and SFT successfully inject the long reasoning data into the VLM models, achieving great improvement on in-domain text-only benchmarks. However, MMLU and MMLU-Pro indicate that SFT greatly weakens performance in general domains. Table 5 shows similar results. In-domain text-only long CoT data can improve a VLM’s performance on in-domain multimodal reasoning. However, results on MMMU and MMMU-Pro indicate that SFT suffers a significant drop in multi-disciplinary multimodal understanding. In contrast, PSFT consistently maintains strong performance, achieving the best overall results across both in-domain and out-of-domain benchmarks.
@@ 5.1 Clipped Tokens in PSFT
Figure 10 presents representative clipped tokens observed during the PSFT training process in Exp. 4.1. The clipped token mostly concentrates on uncertain words like “wait”, “alternatively”, and similar words that reflect certain “long thinking patterns”. With the training in progress, these token-clip weights are more pronounced, while other tokens become smaller. The “thinking pattern” has been gradually and smoothly incorporated into models via PSFT with minimal disturbance to the general capabilities.
@@ 5.2 Parameter Analysis on Clipped Value
As shown in Figure 11, PSFT without clipping maintains higher entropy but suffers from large, unstable gradient norms and highly fluctuating downstream results. By introducing the trust-region clipping mechanism, PSFT strikes a better balance: it prevents entropy collapse, stabilizes gradient updates, and achieves steady improvements in performance. Across different clipping thresholds, with medium values of (e.g., 0.28) showing particularly favorable trade-offs between stability and performance. Flexible choices can be made within an appropriate range of for practical demands.
5.3 Update Frequency of the old policy
| Benchmark | AIME-24 | AIME-25 | TruthfulQA | IFEval |
| Base | 11.25 | 8.75 | 66.10 | 73.94 |
| PSFT (no upd) | 13.02 | 11.15 | 66.61 | 76.46 |
| PSFT (upd.16) | 21.56 | 23.54 | 66.10 | 69.33 |
| PSFT (upd.8) | 19.38 | 21.98 | 67.16 | 73.03 |
| PSFT (upd.4) | 22.50 | 22.92 | 66.72 | 72.08 |
| SFT | 22.08 | 23.02 | 63.14 | 54.42 |
Figure 12 and Table 6 show that dynamically updating the old policy is essential for effective PSFT. While fixed references (no updates) still yield improvements over the base model, their overall performance remains poor. In contrast, frequent updates—especially every 4 steps—deliver substantial gains on in-domain tasks. For update intervals of 16 steps (batch size 256, mini-batch size 16) and 8 steps (mini-batch size 32), we use the same learning rate of . These results further indicate that smaller batch sizes, combined with relatively higher learning rates, can enhance in-domain performance, albeit at the expense of generalization due to high fluctuations.
@@ 6 Related Work
Supervised Fine-tuning. SFT, typically as the first post-training step, sets the stage for a more robust fine-tuned model. Several studies have noted that the standard cross-entropy (CE) loss may not be the most suitable objective for SFT (Li et al., 2024b; Xiao, 2024), arguing that training with CE at this stage tends to encourage the model to memorize the training data rather than acquire more generalizable capabilities. One line of research focuses on improving target performance by addressing the distribution mismatch between expert demonstrations and the evolving target policy. iw-SFT (Qin and Springenberg, 2025) reinterpret standard SFT as optimizing a loose lower bound on the RL objective. This lower bound becomes increasingly loose as the model distribution drifts away from the reference policy. iw-SFT addresses this gap by introducing importance re-weighting, which assigns higher weights to more preferred trajectories and thus tightens the bound toward the true RL objective. DFT (Wu et al., 2025) views SFT as a flawed policy gradient and introduces a simple probability-based reweighting that improves targeted performance. Our work aims to prevent entropy collapse while enhancing target performance, without compromising general capability, and leaving room for further optimization.
@@
Reinforcement Learning. Following SFT, RL is employed to optimize reward signals directly. Policy gradient methods (Sutton et al., 2000) provide the foundation but are prone to instability and high variance. To address this, TRPO (Schulman et al., 2015) introduces a trust-region constraint, limiting each policy update to a small KL-divergence neighborhood to ensure stable improvement. PPO (Schulman et al., 2017) further simplifies this approach with a clipped surrogate objective, striking a balance between stability and efficiency. GRPO (Shao et al., 2024) removes the critic and computes advantage using the relative rewards of outputs within each group, treating all tokens in a sequence equally. We extend the benefits of these methods to the supervised setting.
Combination of SFT and RL. Recent studies have also investigated hybrid objectives that couple SFT loss on offline data with RL loss on online rollouts (Kang et al., 2023; Yan et al., 2025). LUFFY (Yan et al., 2025) exemplifies this direction by mixing offline demonstrations with online trajectories at a fixed proportion in each training batch. Extending this approach, Hybrid Post-Training (HPT) (Lv et al., 2025) adaptively switches between SFT and RL based on rollout performance, balancing guidance and exploration to enhance reasoning. This idea of constructing a composite loss has also been adopted in several recent frameworks (Zhang et al., 2025; Liu et al., 2025). Our method is orthogonal to these methods.
7 Conclusion and Future Work
@@
Motivated by TRPO and PPO, we propose Proximal Supervised Fine-Tuning (PSFT). PSFT preserves a smooth entropy curve while achieving performance comparable to standard SFT, and it exhibits strong generalization. Furthermore, using PSFT as a cold-start model facilitates more effective post-training techniques such as RL optimization and DPO. Overall, PSFT presents a promising alternative to traditional SFT. In the future, we will explore the universality of our PSFT on more diverse, industry-level datasets and models.
Reproducibility statement
Our work is easy to reproduce, as mainstream RL frameworks can support this algorithm with minimal code modifications. You can replace rollout generation with SFT demonstrations and assign a constant value to all advantages in the sequence for training. We provide our code, built upon the open-source verl framework, in the supplementary material. Detailed hyperparameter settings are provided in the Appendix.
Ethics statement
All datasets and models used in this study for our experiments are publicly available and open-source.
Acknowledgement
This work is supported by the AI for Science Program, Shanghai Municipal Commission of Economy and Informatization AI (2025-GZL-RGZN-BTBX-01013), and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
Appendix A Limitation
Our experiments are currently limited in terms of model scale, dataset size, and data diversity. Future work is needed to evaluate the effectiveness of PSFT on industrial-scale models, larger datasets, and mixed data. We have not yet tested our approach on the latest model architectures, such as MoE or hybrid models, which will be considered in future studies.
Appendix B LLM-Usage Statement
We use LLMs to polish writing and correct grammar.
Appendix C Appendix
C.1 Convergence Analysis
Assumption 1 (Smoothness of the NLL)
The negative log likelihood is –smooth, i.e.
Assumption 2 (Bounded stochastic gradients)
There exists such that
Theorem 1 (Convergence Rate )
Under Assumptions 1 and 2, running gradient ascent with learning rate for iterations yields:
Proof. Denote by the objective. Let be the (mini-batch) stochastic gradient at iteration :
Because in the trust region, Assumption 2 implies .
(1) Smoothness of .
For any ,
The same inequality with the opposite sign also holds, so is –smooth.
(2) One-step progress.
With the ascent update and the lower bound of –smoothness,
Taking conditional expectation ,
(3) Bounding the gradient norm.
(4) Choosing the step size.
Choose and assume (so that ). Then
(4) Averaged bound.
When the second term vanishes as , so the dominant rate is . This is consistent with the optimal upper bound of the standard non-convex SGD, indicating that under reasonable assumptions of smoothness and gradient bound, the stochastic gradient rise of PSFT has the same theoretical convergence rate as that of ordinary NLL-SFT.
C.2 Entropy As advantage estimation
We perform training on the high-entropy tokens (Wang et al., 2025).
| Benchmark | AIME-24 | AIME-25 | TruthfulQA |
| PSFT | 26.98 | 28.75 | 62.90 |
| PSFT-Ent( ) | 25.21 | 28.65 | 62.48 |
| PSFT-Ent( ) | 24.17 | 25.10 | 63.71 |
Results.
We believe that this sparse update may also contribute to more robust generalization.
C.3 More large Models
We test PSFT on Qwen3-30B-A3B-2507-Instruct MoE model.
| Benchmark | AIME-24 | AIME-25 | TruthfulQA | IFEval |
| PSFT | 61.67 | 48.38 | 80.19 | 88.52 |
| SFT | 59.89 | 46.88 | 77.89 | 50.56 |
Results.
The results show that PSFT serves as a powerful and practical tool for industrial-scale model training.
C.4 Recent Baselines
Baseline 1: DFT.
DFT (Wu et al., 2025) views SFT as a flawed policy gradient and introduces a simple probability-based reweighting that improves targeted performance.
Baseline 2: Iw-SFT
iw-SFT (Qin and Springenberg, 2025) reframes standard SFT as optimizing a loose lower bound of the RL objective. This bound becomes increasingly loose as the model distribution diverges from the reference policy. To mitigate this mismatch, iw-SFT introduces importance re-weighting, assigning higher weights to more preferred trajectories and thereby tightening the bound toward the true RL objective. In addition, iw-SFT incorporates a clipping range in the importance weights; following the settings in the original paper, we adopt hyperparameter values of 0.2 and 1.8.
| Benchmark | AIME-24 | AIME-25 | TruthfulQA | IFEval |
| Base | 11.25 | 8.75 | 66.10 | 73.94 |
| PSFT (lr 1e-6, bs 32) | 19.38 | 21.98 | 67.16 | 73.03 |
| PSFT (lr 3e-6, bs 32) | 22.50 | 22.92 | 66.63 | 71.32 |
| iw-SFT | 17.64 | 20.31 | 63.67 | 43.92 |
| DFT | 10.73 | 10.10 | 54.53 | 47.34 |
Results.
As shown in Figure 15, DFT exhibits markedly different training dynamics and yields poor performance, which may indicate that it only benefits from domain-specific signals. In contrast, iw-SFT shows training behavior similar to PSFT, though PSFT ultimately achieves better performance.
C.5 Lower learning rate
In this section, we examine the impact of the learning rate on both PSFT and SFT. As shown in Figure 17, a large learning rate induces greater update fluctuations, which in turn results in a higher clip fraction. Moreover, increasing the learning rate enables the model to fit the data distribution more rapidly. Based on Table 10, PSFT still achieves stable performance due to the gradient clipping mechanism, and the evaluation on IFEval shows that PSFT outperforms SFT with a small learning rate.
| Benchmark | AIME-24 | AIME-25 | TruthfulQA | IFEval |
| Base | 11.25 | 8.75 | 66.10 | 73.94 |
| PSFT (lr 1e-6, bs 32) | 19.38 | 21.98 | 67.16 | 73.03 |
| PSFT (lr 3e-6, bs 32) | 22.50 | 22.92 | 66.63 | 71.32 |
| SFT (lr 2e-5, bs 256) | 22.08 | 23.02 | 63.14 | 54.42 |
| SFT (lr 5e-6, bs 256) | 21.25 | 22.50 | 66.54 | 68.87 |
C.6 Coding Task
We use the OpenR1-code dataset and filter it for lengths less than 8192, composing the 13k dataset. We use LiveCodeBench as the evaluation (Jain et al., 2025).
Results.
These results show that while both SFT and PSFT effectively learn the target domain, PSFT better preserves mathematical instruction-following ability.
C.7 Experimental Details